A metal wire 75.0 cm long and 0.130 cm in diameter stretches 0.0350 cm when a load of 8.00 kg is hung on its end. Find the stress, the strain, and the Young’s modulus for the material of the wire.


1
Expert's answer
2022-02-07T15:27:23-0500

The stress, the strain, and the Young’s modulus for the material of the wire respectively are:


"\sigma=\dfrac {4mg}{\pi d^2}==59·10^6\text{ Pa}.\\\space\\
\epsilon=\dfrac{\Delta L}L·100\%=0.047\%.\\\space\\
Y=\dfrac\sigma\epsilon=1.26·10^{11}\text{ Pa}."